OpenStudy (immings):

Convert 0.575% (w/w) CuSO4 5H2O to %(w/v) Density of 0.575%(w/w) is 1.67g/mL Molar mass = 249.7g/mol

OpenStudy (anonymous):

we have to find 1st of all the volume. 0.575% (w/w) meant 0.575 g of Cuso4.5h2O are dissolved in 100 gram of sample. density = mass / volume OR volume = mass / density so, vol = 100 / 1.67 = 59.88 mL its mean 0.575 g of solute is dissolved in 59.88 mL, so %age will be: w/V% = 0.575 / 59.88 X 100 = 0.96 % (w/V)